Compute W(y1,y2) where y1 and y2 are solutions of the differential equation y"+y'+y=0


1
Expert's answer
2021-01-06T19:58:35-0500


y''+y'+y=0 is a linear homogeneous differential equation with constant coefficients.

The solution is sought in the form of an exponent y = "e^{\lambda x}" .

We write the characteristic equation assuming that

y' = "\lambda""e^{\lambda x}" ; y'' = "\lambda"2"e^{\lambda x}"

Hence

"\lambda"2"e^{\lambda x}" + "\lambda""e^{\lambda x}" + "e^{\lambda x}" = 0 or "e^{\lambda x}" ( "\lambda"2 + "\lambda" + 1) = 0

"\lambda"2 + "\lambda" + 1 = 0 is the caracteristic equation.

The above equation is the standard quadratic equation:

D = b2 - 4ac; a=b=c=1

D = 12 -4*1*1 = -3

"\lambda"1,2 = "\frac{ - b ± \sqrt{\smash[b]{D}}}{2a}"

"\lambda"1 = "\frac{ - 1 + i \sqrt{\smash[b]{3}}}{2}"

"\lambda"2 = "\frac{ - 1 - i \sqrt{\smash[b]{3}}}{2}"

y = c1exp( "\lambda"1x) + c2exp( "\lambda"2x)

y1 = c1exp( "\frac{ - 1 + i \sqrt{\smash[b]{3}}}{2}"x)

y2 = c2exp( "\frac{ - 1 - i \sqrt{\smash[b]{3}}}{2}"x)

If y = u(x)+iv(x) is a solution to the equation then y1=u(x) and y2=iv(x) (here i = "\sqrt-1" ) are solutions too.

Hence:

we can write

y1=c1"e^\frac{-x}{2}"sin("\frac{\sqrt3}{2}"x);

y2=c2"e^\frac{-x}{2}" cos("\frac{\sqrt3}{2}"x)

W = "\begin{vmatrix}
 y_1 & y_2 \\
 y'_1 & y'_2
\end{vmatrix}"

y'1 = c1"\lbrace" -"\frac{1}{2}""e^\frac{-x}{2}"sin("\frac{\sqrt{\smash[b]{3}}}{2}"x) +"\frac{\sqrt{\smash[b]{3}}}{2}""e^\frac{-x}{2}"cos("\frac{\sqrt{\smash[b]{3}}}{2}"x)"\rbrace"

y'2 = c2"\lbrace" -"\frac{1}{2}""e^\frac{-x}{2}"cos("\frac{\sqrt{\smash[b]{3}}}{2}"x) -"\frac{\sqrt{\smash[b]{3}}}{2}""e^\frac{-x}{2}"sin("\frac{\sqrt{\smash[b]{3}}}{2}"x)"\rbrace"


W = "\begin{vmatrix}
 c_1e^\frac{-x}{2}sin(\frac{\sqrt3}{2}x) & c_2e^\frac{-x}{2}cos(\frac{\sqrt3}{2}x) \\
 \frac{c_1e^\frac{-x}{2}}{2}\lbrace-sin(\frac{\sqrt3}{2}x)+\sqrt3cos(\frac{\sqrt3}{2}x)\rbrace & \frac{-c_2e^\frac{-x}{2}}{2}\lbrace cos(\frac{\sqrt3}{2}x)+\sqrt3sin(\frac{\sqrt3}{2}x)\rbrace
\end{vmatrix}" =

= "c_1e^\frac{-x}{2}sin(\frac{\sqrt3}{2}x)(\frac{-c_2e^\frac{-x}{2}}{2})\lbrace cos(\frac{\sqrt3}{2}x)+\sqrt3sin(\frac{\sqrt3}{2}x)\rbrace-"

"- c_2e^\frac{-x}{2}cos(\frac{\sqrt3}{2}x)(\frac{c_1e^\frac{-x}{2}}{2})\lbrace-sin(\frac{\sqrt3}{2}x)+\sqrt3cos(\frac{\sqrt3}{2}x)\rbrace ="


"= - \frac{c_1c_2}{2}e^{-x}\lbrace \bcancel{cos(\frac{\sqrt3}{2}x)sin(\frac{\sqrt3}{2}x)} - \bcancel{cos(\frac{\sqrt3}{2}x)sin(\frac{\sqrt3}{2}x)} +
 \\ +\sqrt3sin^2(\frac{\sqrt3}{2}x)+\sqrt3cos^2(\frac{\sqrt3}{2}x) \rbrace="

"=- \frac{\sqrt3}{2}c_1c_2e^{-x}"

"if\\ c_1=c_2=1 \\W = =- \frac{\sqrt3}{2}e^{-x}"